Consider an infinite slab of thickness d and uniform volume charge density ρ. This is essentially an infinite plane with a non-negligible thickness. The planar symmetry involves z to -z reflection symmetry, as well as infinite in the x-y plane. Use Gauss law to find the electric field produced outside the slab and inside the slab .


1
Expert's answer
2021-10-26T17:20:38-0400

Electric flux "\phi=\frac{q}{\epsilon_0}"

"\phi=EAcos\theta"


"\rho=\frac{q}{v}"

"\Phi=\smallint E.ds_1+\smallint E.ds_2+\int E.ds_3"

"\Phi=ES+ES+0=2ES"

"2ES=\frac{\rho S}{2\epsilon}"

"E=\frac{\rho }{2\epsilon_0 }"